# Risk-Averse Learning with Delayed feedback

### v1

@@ 3.1 One-point risk-averse learning
    when $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}$ and $\eta=(T+D)^{-\frac{3}{4}}$, Algorithm 1 achieves ${\rm R}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\Big{)}$ with high probability.
Proof. Denoting $x_{\delta}^{\ast}=\operatorname*{arg\,min}_{x\in\mathcal{X}^{\delta}}\sum_{t=1}^{T}C_{t}^{\delta}(x)$, we have
$\sum_{t=1}^{T}C_{t}^{\delta}(x_{\delta}^{\ast})$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}C_{t}^{\delta}((1-\delta/r)x)$
$\leq\min_{x\in\mathcal{X}}\sum_{t=1}^{T}(\delta/r)C_{t}^{\delta}(0)+(1-\delta/r)C_{t}^{\delta}(x)$
$\leq\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\Big{(}C_{t}^{\delta}(x)+(\delta/r)L_{0}\left\|x\right\|\Big{)}$
$\leq D_{x}L_{0}T\delta/r+\min_{x\in\mathcal{X}}\sum_{t=1}^{T}C_{t}^{\delta}(x),$ (10)
where the first inequality is from the convexity of $C_{t}^{\delta}$ as in Lemma 1, and the second inequality is from Lipschitzness of $C_{t}^{\delta}$ as in Lemma 2. The regret (1) is written as
$\mathrm{R}_{1}(T)$ $\leq\sum_{t=1}^{T}C_{t}^{\delta}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x^{\ast})+2\delta L_{0}T$
$\leq\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x^{\ast})+3\delta L_{0}T$
$\leq\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{\delta}^{\ast})+(3+D_{x}/r)\delta L_{0}T,$ (11)
@@
where the first inequality is from the definition of $C_{t}^{\delta}$, as in (2), the second inequality is from the Lipschitzness of $C_{t}^{\delta}$, and the third inequality is from substituting (3.1) into (3.1). Denoting $\tilde{C}_{\tau}(\cdot)=C_{\mu(\tau)}^{\delta}$, we have
$\hskip 10.00002pt\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{\delta}^{\ast})$
$\leq\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)})-\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\delta}^{\ast})+2U\bar{d}$
$=\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau-\tilde{s}_{\tau}})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(x_{\delta}^{\ast})+2U\bar{d}$
$=\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau-\tilde{s}_{\tau}})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})$
$\hskip 10.00002pt+\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(x_{\delta}^{\ast})+2U\bar{d},$ (12)
where the first inequality follows from $T-L_{T}\leq\bar{d}$ and $|C_{t}(\cdot)|\leq U$, for all $t$. The first equality follows from $x_{t}=\tilde{x}_{L_{t-1}+1}$ and $x_{\mu(\tau)}(\cdot)=\tilde{x}_{L_{\mu(\tau)-1}+1}=\tilde{x}_{\tau-\tilde{s}_{\tau}}$. Denote $\tilde{g}_{\tau}=g_{\mu(\tau)}=g_{\mu(\tau)|\mu(\tau)+d_{\mu(\tau)}}$, which represents the estimated gradient (9) at time slot $t=\mu(\tau)+d_{\mu(\tau)}$. Since $\|\tilde{x}_{\tau}-\tilde{x}_{\tau+1}\|=\|\tilde{x}_{\tau}-\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{\tau}-\eta\tilde{g}_{\tau})\|\leq\eta\|\tilde{g}_{\tau}\|\leq\frac{d\eta}{\delta}U$, we have
$\mathbb{E}[\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|]$ $\leq\mathbb{E}[\sum_{j=0}^{\tilde{s}_{\tau}-1}\|\tilde{x}_{\tau-\tilde{s}_{\tau}+j}-\tilde{x}_{\tau-\tilde{s}_{\tau}+j+1}\|]$
$\leq\mathbb{E}[\eta\tilde{s}_{\tau}\|\tilde{g}_{\tau}\|]\leq\frac{d\eta}{\delta}U\tilde{s}_{\tau}.$ (13)
For the first two terms of (3.1), we have
$\hskip 10.00002pt\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau-\tilde{s}_{\tau}})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})\leq\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla\tilde{C}_{\tau},\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\big{\rangle}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}L_{0}\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|\Big{]}\leq\sum_{\tau=1}^{L_{T}}\frac{\eta L_{0}dU}{\delta}\tilde{s}_{\tau}=\frac{\eta L_{0}dU}{\delta}D,$ (14)
@@
where the first inequality is from the convexity of $C_{t}(\cdot)$ and the second inequality is from the Lipschitzness of $C_{t}(\cdot)$. The third inequality is from substituting (3.1) into (3.1). The equality comes from the lemma 4. By the update rule (9), we have
$\hskip 13.00005pt\|\tilde{x}_{\tau+1}-x_{\delta}^{\ast}\|^{2}$
$=\|\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{\tau}-\eta\tilde{g}_{\tau})-x_{\delta}^{\ast}\|^{2}$
$\leq\|\tilde{x}_{\tau}-\eta\tilde{g}_{\tau}-x_{\delta}^{\ast}\|^{2}$
$=\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|^{2}+\eta^{2}\|\tilde{g}_{\tau}\|^{2}-2\eta\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-x_{\delta}^{\ast}\rangle,$
@@
where the inequality follows from the fact that $x_{\delta}^{\ast}\in\mathcal{X}^{\delta}$. Then, we obtain
$\hskip 10.00002pt\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-x_{\delta}^{\ast}\rangle$
$\leq\frac{\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|^{2}-\|\tilde{x}_{\tau+1}-x_{\delta}^{\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}.$ (15)
@@
Note that constructing the empirical distribution function of the cost function using finite samples induces CVaR gradient estimate error:
$\hat{\varepsilon}_{s_{n}|t}:={\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}]-{\rm CVaR}_{\alpha}[F_{s_{n}|t}],$ (16)
for $n=1,\dots,\Delta_{t}$ and $t=1,\dots,T$. For the third and fourth terms of (3.1), we have
$\hskip 10.00002pt\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(\tilde{x}_{\tau})-\sum_{\tau=1}^{L_{T}}\tilde{C}_{\tau}(x_{\delta}^{\ast})$
$\leq\sum_{\tau=1}^{L_{T}}\left\langle\nabla\tilde{C}_{\tau},\tilde{x}_{\tau}-x_{\delta}^{\ast}\right\rangle$
$=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\big{\langle}\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau}-x_{\delta}^{\ast}\big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\frac{\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|^{2}-\|\tilde{x}_{\tau+1}-x_{\delta}^{\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}\Big{]}$
$\hskip 10.00002pt+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|\Big{]}$
$\leq\frac{D_{x}^{2}}{\eta}+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta^{2}}L_{T}+R_{11},$ (17)
@@
with $R_{11}=\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-x_{\delta}^{\ast}\|]$ with $\tilde{\varepsilon}_{\tau}=\varepsilon_{\mu(\tau)}$. The first inequality follows from the convexity of $C_{t}$, and the first equality follows from substituting (3) and (16) into (3.1). The second inequality follows from substituting (3.1) into (3.1) and the last inequality follows from $D_{x}$ is the diameter of $\mathcal{X}$ and $\mathbb{E}[\|\tilde{g}_{\tau}\|^{2}]\leq\frac{d^{2}U^{2}}{\delta^{2}}$. Leveraging the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [21], we have
@@
with $m_{t}$ being the sampling number at time $t$. Denote the event in (18) as $A_{t}$, and denote $\mathbb{P}\{A_{t}\}$ as the occurrence probability of event $A_{t}$, for $t=1,\dots,L_{T}$. By Lemma 3, the estimation error (16) is bounded with the difference between the true and the empirical probability function:
$|\hat{\varepsilon}_{t}|=\frac{U}{\alpha}\sup\big{|}\hat{F}_{t}-F_{t}\big{|}\leq\frac{U}{\alpha}\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}},$ (19)
@@
with probability at least $1-\bar{\gamma}$, for $t=1,\dots,L_{T}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, by substituting (19) into $R_{11}$, we obtain that
@@
$R_{11}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\tilde{\varepsilon}_{\tau}\|]$
$\leq\frac{dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2m_{\tau}}}$
$\leq\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2}}L_{T}^{1-\frac{a}{2}},$ (20)
@@
with probability at least $1-\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{L_{T}}A_{\tau}\}\geq 1-\sum_{\tau=1}^{L_{T}}\mathbb{P}\{A_{\tau}\}=1-L_{T}\frac{\gamma}{L_{T}}=1-\gamma$. Substituting (3.1), (3.1), (3.1) and (3.1) into (3.1), with $L_{T}\leq T$, we have
${\rm R}_{1}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+\frac{\eta L_{0}dU}{\delta}D+\frac{D_{x}^{2}}{\eta}$
$+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta^{2}}T+\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2}}T^{1-\frac{a}{2}},$ (21)

### v2

@@ 3.1 One-point risk-averse learning
    when $a\geq 1$, selecting $\delta=T^{-\frac{1}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3}{4}}\sqrt{\bar{d}P_{T}}$, Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}(T+D)^{\frac{3}{4}}\sqrt{\bar{d}P_{T}}\Big{)}$ with high probability.
Proof. The regret (2) is written as
$\hskip-5.0pt\mathrm{R}_{1}(T)$ $=\sum_{t=1}^{T}\mathbb{E}_{u}[C_{t}(\hat{x}_{t}+\delta u-\delta u)]-\mathbb{E}_{u}[C_{t}(x_{t}^{\ast}+\delta u-\delta u)]$
$\leq\sum_{t=1}^{T}\Big{(}C_{t}^{\delta}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\ast})\Big{)}+2\delta L_{0}T$
$\leq\sum_{t=1}^{T}\Big{(}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\ast})\Big{)}+3\delta L_{0}T,$ (13)
@@
where the first inequality is from the Lipschitzness of $C_{t}$ and the definition of $C_{t}^{\delta}$, as in (3). The second inequality is from the Lipschitzness of $C_{t}^{\delta}$. Furthermore, for $t=1,\dots,T$, we have
$\min_{x_{t}\in\mathcal{X}^{\delta}}C_{t}^{\delta}(x_{t})$ $=\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}((1-\delta/r)x_{t})$
$\leq\min_{x_{t}\in\mathcal{X}}(\delta/r)C_{t}^{\delta}(0)+(1-\delta/r)C_{t}^{\delta}(x_{t})$
$\leq\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}(x_{t})+(\delta/r)L_{0}\left\|x_{t}\right\|$
$\leq\min_{x_{t}\in\mathcal{X}}C_{t}^{\delta}(x_{t})+D_{x}L_{0}\delta/r,$ (14)
@@
where the first inequality is from the convexity of $C_{t}^{\delta}$, as in Lemma 1, and the second inequality is from Lipschitzness of $C_{t}^{\delta}$, as in Lemma 2. Denote $x_{t}^{\delta,\ast}={\textrm{argmin}}_{x\in\mathcal{X}^{\delta}}C_{t}^{\delta}(x)$. Substituting (3.1) into (3.1), we have
$\mathrm{R}_{1}(T)\leq\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\delta,\ast})+(3+D_{x}/r)\delta L_{0}T,$ (15)
Denote $\tilde{C}_{\tau}(\cdot)=C_{\mu(\tau)}^{\delta}$. The first and second terms of (15) is further written as
$\hskip 10.00002pt\sum_{t=1}^{T}C_{t}^{\delta}(x_{t})-\sum_{t=1}^{T}C_{t}^{\delta}(x_{t}^{\delta,\ast})$
$\leq\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)})-\sum_{\tau=1}^{L_{T}}C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}^{\delta,\ast})+2U\bar{d}$
$\leq\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}+2U\bar{d}$
$=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-\tilde{x}_{\tau}+\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}$
$\hskip 10.00002pt+\tilde{x}_{\tau}^{\delta,\ast}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}+2U\bar{d},$
$=R_{1}+R_{2}+R_{3}+2U\bar{d}$ (16)
with $R_{1}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),x_{\mu(\tau)}-\tilde{x}_{\tau}\big{\rangle}$, $R_{2}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\big{\rangle}$ and $R_{3}=\sum_{\tau=1}^{L_{T}}\big{\langle}\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)}),\tilde{x}_{\tau}^{\delta,\ast}-x_{\mu(\tau)}^{\delta,\ast}\big{\rangle}$. In (3.1), the first inequality follows from $T-L_{T}\leq\bar{d}$ and $|C_{t}(\cdot)|\leq U$. The second inequality is from the convexity of $C_{t}$, for all $t$. In the following, we are going to bound $R_{1}$, $R_{2}$ and $R_{3}$.
Denote $\tilde{g}_{\tau}$ as the $\tau$-th CVaR gradient estimate being used for gradient descent (11), for $\tau=1,\dots,L_{T}$. Since $x_{t}=\tilde{x}_{L_{t-1}+1}$ and $x_{\mu(\tau)}=\tilde{x}_{L_{\mu(\tau)-1}+1}=\tilde{x}_{\tau-\tilde{s}_{\tau}}$, we have
$\hskip 10.00002pt\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|$
$=\|\tilde{x}_{\tau-\tilde{s}_{\tau}}-\tilde{x}_{\tau}\|\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\|\tilde{x}_{\tau-\tilde{s}_{\tau}+j}-\tilde{x}_{\tau-\tilde{s}_{\tau}+j+1}\|$
$\leq\sum_{j=0}^{\tilde{s}_{\tau}-1}\eta\|\tilde{g}_{\tau-\tilde{s}_{\tau}+j}\|\leq\frac{d\eta}{\delta}U\tilde{s}_{\tau},$ (17)
where the second inequality is from $\|\tilde{x}_{i}-\tilde{x}_{i+1}\|=\|\tilde{x}_{i}-\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{i}-\eta\tilde{g}_{i})\|\leq\eta\|\tilde{g}_{i}\|$ and the last inequality is from $\|\tilde{g}_{i}\|\leq\frac{dU}{\delta}$, for $i\in\{\tau-\tilde{s}_{\tau},\dots,\tau-1\}$, according to its definition specified in (10). Thus, for $R_{1}$ in (3.1), we have
$R_{1}$ $\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}L_{0}\|x_{\mu(\tau)}-\tilde{x}_{\tau}\|\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\frac{\eta L_{0}dU}{\delta}\tilde{s}_{\tau}=\frac{\eta L_{0}dUD}{\delta},$ (18)
where the first inequality is Lemma 2, the second inequality is from substituting (3.1) into (3.1) and the equality is from the definition of cumulative delay as in (12).
The next is to bound $R_{2}$ in (3.1). The equality is from Lemma 4. By the gradient descent (11), we have
$\hskip 13.00005pt\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$=\|\mathcal{P}_{\mathcal{X}^{\delta}}(\tilde{x}_{\tau}-\eta\tilde{g}_{\tau})-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$\leq\|\tilde{x}_{\tau}-\eta\tilde{g}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}$
$=\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}+\eta^{2}\|\tilde{g}_{\tau}\|^{2}-2\eta\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\rangle,$
@@
where the inequality follows from the fact that $\tilde{x}_{\tau}^{\delta,\ast}\in\mathcal{X}^{\delta}$. Then, we obtain
$\hskip 10.00002pt\langle\tilde{g}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\rangle$
$\leq\frac{\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}.$ (19)
@@
Note that constructing the empirical distribution function of the cost function using finite samples induces CVaR estimate error $\hat{\epsilon}_{t}:={\mathrm{CVaR}}_{\alpha}[\hat{F}_{t}]-{\mathrm{CVaR}}_{\alpha}[F_{t}]$, where $\hat{F}_{t}$ is the empirical distribution function that corresponds to $F_{t}$, for $t=1,\dots,T$. From the gradient of smoothed function (4) and the definition of $\hat{g}_{t}^{n}$ given in (10), the error of $\tau$-th utilized CVaR gradient estimate is
$\nabla C_{\mu(\tau)}^{\delta}(x_{\mu(\tau)})=\mathbb{E}[\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\epsilon}_{\tau}\tilde{u}_{\tau}].$ (20)
for $\tau=1,\dots,L_{T}$. Then, for $R_{2}$, we have
$R_{2}$ $=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\big{\langle}\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\frac{\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}}{2\eta}+\frac{\eta}{2}\|\tilde{g}_{\tau}\|^{2}\Big{]}$
$\hskip 10.00002pt+\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|\Big{]}$
$\leq R_{21}+R_{22}+\frac{\eta}{2}\frac{d^{2}U^{2}}{\delta^{2}}L_{T},$ (21)
@@
where $R_{21}=\frac{1}{{2\eta}}\sum_{\tau=1}^{L_{T}}\big{(}\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}-\|\tilde{x}_{\tau+1}-\tilde{x}_{\tau}^{\delta,\ast}\|^{2}\big{)}$ and $R_{22}=\sum_{\tau=1}^{L_{T}}\mathbb{E}[\|\frac{d}{\delta}{\tilde{\varepsilon}_{\tau}}\tilde{u}_{\tau}\|\|\tilde{x}_{\tau}-\tilde{x}_{\tau}^{\delta,\ast}\|]$. The first equality follows from substituting (20) into (3.1). The first inequality follows from substituting (3.1) into (3.1) and the second inequality follows from $\mathbb{E}[\|\tilde{g}_{\tau}\|^{2}]\leq\frac{d^{2}U^{2}}{\delta^{2}}$. Furthermore, we have
$R_{21}$ $=\frac{1}{2\eta}\sum_{\tau=1}^{L_{T}}\left(\|\tilde{x}_{\tau}\|^{2}-\|\tilde{x}_{\tau+1}\|^{2}+2(\tilde{x}_{\tau}^{\delta,\ast})^{\top}(\tilde{x}_{\tau+1}-\tilde{x}_{\tau})\right)$
$\leq\frac{1}{2\eta}\Big{(}D_{x}^{2}+2\sum_{t=2}^{L_{T}}\tilde{x}_{\tau}^{\top}(\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast})$
$\hskip 10.00002pt+2(\tilde{x}_{L_{T}}^{\delta,\ast})^{\top}\tilde{x}_{L_{T}+1}-2(\tilde{x}_{1}^{\delta,\ast})^{\top}\tilde{x}_{2}\Big{)}$
$\leq\frac{1}{2\eta}\left(3D_{x}^{2}+2D_{x}\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast}\|\right)$
$\leq\frac{1}{2\eta}\left(3D_{x}^{2}+2D_{x}\sum_{\tau=2}^{L_{T}}\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|\right)$
$=\frac{1}{2\eta}\left(3D_{x}^{2}+4D_{x}\bar{d}P_{T}\right)$ (22)
where the first and the second inequalities are from the fact that the diameter of $\mathcal{X}$ is $D_{x}$. The third inequality is from the projection mapping and $\|\tilde{x}_{\tau-1}^{\delta,\ast}-\tilde{x}_{\tau}^{\delta,\ast}\|=(1-\delta/r)\|\tilde{x}_{\tau-1}^{\ast}-\tilde{x}_{\tau}^{\ast}\|$ and the last equality is from Lemma 5.
The next is to bound $R_{22}$. Leveraging the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality [26], we have
@@
with $m_{t}$ being the sampling number at time $t$. Denote the event in (23) as $A_{t}$, and denote $\mathbb{P}\{A_{t}\}$ as the occurrence probability of event $A_{t}$, for $t=1,\dots,L_{T}$. By Lemma 3, the estimation error (20) is bounded by
$|\hat{\varepsilon}_{t}|=\frac{U}{\alpha}\sup\big{|}\hat{F}_{t}-F_{t}\big{|}\leq\frac{U}{\alpha}\sqrt{\frac{\ln(2/\bar{\gamma})}{2m_{t}}}$ (24)
@@
with probability at least $1-\bar{\gamma}$, for $t=1,\dots,L_{T}$. Let $\gamma=\bar{\gamma}L_{T}$. Then, by substituting (24) into $R_{22}$, we obtain that
@@
$R_{22}$ $\leq\frac{dD_{x}}{\delta}\sum_{\tau=1}^{L_{T}}\mathbb{E}[|\tilde{\varepsilon}_{\tau}|]\leq\frac{dUD_{x}}{\alpha\delta}\sum_{\tau=1}^{L_{T}}\sqrt{\frac{\ln(2L_{T}/\gamma)}{2\tilde{m}_{\tau}}}$
$\leq\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (25)
@@
with probability at least $1-\gamma$, which establishes as $1-\mathbb{P}\{\bigcup_{\tau=1}^{L_{T}}A_{\tau}\}\geq 1-\sum_{\tau=1}^{L_{T}}\mathbb{P}\{A_{\tau}\}=1-L_{T}\frac{\gamma}{L_{T}}=1-\gamma$ and $L_{T}\leq T$.
For $R_{3}$ in (3.1), we have
$R_{3}$ $=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\Big{\langle}\tilde{g}_{\tau}-\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau},\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\Big{\rangle}\Big{]}$
$\leq\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\Big{(}\|\tilde{g}_{\tau}\|+\Big{\|}\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau}\Big{\|}\Big{)}\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$
$\leq R_{31}+R_{32}$ (26)
with $R_{31}=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\tilde{g}_{\tau}\|\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$ and $R_{32}=\sum_{\tau=1}^{L_{T}}\mathbb{E}\Big{[}\|\frac{d}{\delta}\tilde{\varepsilon}_{\tau}\tilde{u}_{\tau}\|\big{\|}\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\big{\|}\Big{]}$. The first equality follows from substituting (20) into (3.1). Similar to (3.1), we have
$R_{32}\leq\frac{cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}}.$ (27)
Furthermore, for $R_{31}$, we have
$R_{31}$ $\leq\frac{dU}{\delta}\sum_{\tau=1}^{L_{T}}\|\tilde{x}_{\tau}^{\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\ast}\|\leq\frac{2dU}{\delta}\bar{d}P_{T}$ (28)
where the first inequality is from $\|\tilde{g}_{\tau}\|\leq\frac{dU}{\delta}$ and $\|\tilde{x}_{\tau}^{\delta,\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\delta,\ast}\|=(1-\delta/r)\|\tilde{x}_{\tau}^{\ast}-\tilde{x}_{\tau-\tilde{s}_{\tau}}^{\ast}\|$. The second inequality is from Lemma 5.
Substituting (3.1), (3.1), (3.1), (3.1), (3.1), (27) and (28) into (15), with $L_{T}\leq T$, we have
${\rm R}_{1}(T)$ $\leq(3+D_{x}/r)\delta L_{0}T+2U\bar{d}+\frac{\eta L_{0}dU}{\delta}D+\frac{\eta d^{2}U^{2}}{2\delta^{2}}T$
$\hskip-30.00005pt+\hskip-3.00003pt\frac{3D_{x}^{2}+4D_{x}\bar{d}P_{T}}{2\eta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2dU\bar{d}P_{T}}{\delta}\hskip-1.99997pt+\hskip-1.99997pt\frac{2cdUD_{x}}{\alpha\delta}\sqrt{\frac{\ln(2T/\gamma)}{2}}T^{1-\frac{a}{2}},$ (29)